OpenStudy (beccab003):

Solve for x in the equation: 2y/3x = 6 Please explain answer. Thanks

OpenStudy (anonymous):

you need to divide both sides of the equation by 2y. To isolate the x. Then you need to divide all of it by three. You'll get a fraction with a y variable.

OpenStudy (luigi0210):

Or multiply both sides by 3x then divide it.

OpenStudy (anonymous):

\[\frac{ 2y }{ 3x } = 6\] \[3x(\frac{ 2y }{ 3x }) = 6 ( 3x)\] \[2y=18x\] \[\frac{ 2y }{ 2 } = \frac{ 18x }{ 2 }\] \[y = 9x\] \[\frac{ y }{ 9 }=\frac{ 9x }{ 9 }\] \[x = \frac{ y }{ 9 }\]

OpenStudy (beccab003):

So would it be: 1) x=1/9y 2) x=9y 3) x=1/3y

OpenStudy (anonymous):

Is this the original equation? \[\frac{ 2y }{ 3x }=6\]

OpenStudy (anonymous):

@BeccaB003 If that is actually the original equation, the correct answer choice is not among those you listed. Can you verify the original question?

OpenStudy (beccab003):

Yes

OpenStudy (beccab003):

This was the original equation: Solve for x in the equation: 2y/3x = 6

OpenStudy (beccab003):

Here are the answer options to the equation: 1) x=1/9y 2) x=9y 3) x=1/3y 4) Impossible to do.

OpenStudy (anonymous):

The correct solution for X is\[x=\frac{ y }{ 9 }\]

OpenStudy (beccab003):

Then what do I do?

OpenStudy (anonymous):

Here. We can CHECK the answer I gave to verify that it is actually correct:

OpenStudy (anonymous):

\[\frac{ 2y }{ 3x } = 6\] \[\frac{ 2y }{ 3(\frac{ y }{ 9 }) }=6\] \[\frac{ 2y }{ (\frac{ 3y }{ 9 }) }=6\] \[\frac{ 18y }{ 3y }=6\] \[6 = 6\] TRUE. Therefore, the correct solution for X is: \[x=\frac{ y }{ 9 }\]

OpenStudy (anonymous):

If this is an online quiz of some sort, then you'll just have to pick something and then let the instructor know that none of the answer choices is correct and explain why, which I have done for you.

geerky42 (geerky42):

I think by 1/9y (first option), he means \(\dfrac{1}{9}y\)

OpenStudy (anonymous):

Ah. Then, that would be the correct answer. Good call, @geerky42

OpenStudy (anonymous):

@BeccaB003 , is that how the answer choice appears? \[\frac{ 1 }{ 9 }y\]

OpenStudy (anonymous):

@BeccaB003 or is it \[\frac{ 1 }{ 9y }\] ?

OpenStudy (beccab003):

Yes. it looks like: \[\frac{ 1 }{ 9 }y\]

OpenStudy (beccab003):

So will that be the correct answer?

OpenStudy (anonymous):

Yup. That is correct.

OpenStudy (beccab003):

Okay great, thanks. What about this problem...It is similar to the other one: Solve for x in the equation: 3y/7x = 9 I think the answer is \[\frac{ 1 }{ 21 }y\]

OpenStudy (beccab003):

In problems like this do you times the equation by the denominator?

OpenStudy (anonymous):

Your ultimate goal in solving any equation with a variable(s) is to isolate the variable or get it by itself on one side of the equation and get a value on the other side of the equation.

OpenStudy (anonymous):

And, yes, your answer is right.

OpenStudy (beccab003):

Great!! Thanks so much. You have been the most helpful person to answer any of my questions. Is there any way for me to ask you more different questions later?

OpenStudy (beccab003):

Like will I still be able to type things to you if I close my current question?

OpenStudy (anonymous):

My pleasure. You know, I'm kind of new to using this site. But I think if you just send me a message with the Mail/Envelope icon at the top of the screen, I will get it. Just type my username, TOMWALK, into the To: field.

OpenStudy (beccab003):

Okay thanks!! I'm new to the site too and I really appreciate your help because I've been having a hard time with math and I have an exam next Friday. Are you a high school student or older and out of school?

OpenStudy (anonymous):

I'm older. I actually tutor middle school/junior high math.

OpenStudy (anonymous):

I tutor in the NW Indiana/Chicagoland area.

OpenStudy (beccab003):

That's cool! Thanks so much for all your help!! I like that your answers explained step by step because I want to learn how to do the problems by myself and if I know how to work them then it is really helpful. Thanks!!!

OpenStudy (anonymous):

I'm happy to help. Just send me a message if I can help with anything else. That's probably easiest.

OpenStudy (beccab003):

Okay thanks! Bye!!